### v1

# IC/DC: Surpassing Heuristic Solvers in Combinatorial Optimization with Diffusion Models

Abstract
Recent advancements in learning-based combinatorial optimization (CO) methods have shown promising results in solving NP-hard problems without the need for expert-crafted heuristics. However, high performance of these approaches often rely on problem-specific human-expertise-based search after generating candidate solutions, limiting their applicability to commonly solved CO problems such as Travelling Salesman Problem (TSP). In this paper, we present IC/DC, a CO framework that operates without any supervision. IC/DC is specialized in addressing problems involving two distinct sets of items, and it does not need problem-specific search processes to generate valid solutions. IC/DC employs a novel architecture capable of capturing the intricate relationships between items, and thereby enabling effective optimization in challenging CO scenarios. We train our model in a self-supervised way to minimize the cost of the solution while adhering to the problem-specific constraints. IC/DC not only achieves state-of-the-art performance compared to previous learning methods, but also surpasses well-known solvers and heuristic approaches on Asymmetric Traveling Salesman Problem (ATSP).
1 Introduction
Combinatorial optimization (CO) aims to find the optimal solution that maximizes or minimizes an objective function from a large, discrete set of feasible solutions. This field has been extensively studied due to its broad industrial applications, including logistics, supply chain optimization, job allocation, and more (Zhang et al. 2023). Despite its significance, many CO problems are NP-complete, and developing efficient approximation algorithms is essential.
Traditionally, approximation algorithms for CO have been developed using mathematical programming or hand-crafted heuristics (Miller, Tucker, and Zemlin 1960; Helsgaun 2023). However, the need for problem-specific expertise and the high computational demands of these methods has sparked increasing interest in applying deep learning techniques to CO problems. Early deep learning approaches framed CO problems as sequential decision-making tasks, generating solutions in an auto-regressive manner (Bello et al. 2016; Kool, Van Hoof, and Welling 2018). However, these methods were relatively limited in performance due to their inability to revise previously made decisions.
In contrast, diffusion-based methods generate complete solutions in a single diffusion timestep and then iteratively refine them during the denoising process (Sun and Yang 2023; Min, Bai, and Gomes 2024). This iterative refinement enhances the overall quality of the solutions through corrections and adjustments. By allowing for the revision of earlier decisions, diffusion-based methods overcome the limitations of auto-regressive approaches and avoid the compounding errors typically associated with early decisions.
Despite the impressive performance of diffusion-based methods, previously proposed algorithms had several significant drawbacks, such as the need for costly supervision (Sun and Yang 2023), or use of problem-specific objective that are applicable to other CO problems (Min, Bai, and Gomes 2024). Additionally, due to the challenges in imposing constraint within a diffusion model, previous studies relied on problem-specific search process to extract feasible solutions from the diffusion model’s generations. Designing these problem-specific search requires specialized knowledge and cannot be easily adapted to different CO problems.
In this work, we propose a novel method for training a diffusion model in a self-supervised manner, eliminating the need for costly supervision, problem-specific objectives, or problem-specific search process. We demonstrate our approach on two distinct and challenging CO problems—the parallel machine scheduling problem (PMSP) and the asymmetric travelling salesman problem (ATSP)—which have received less attention (Kwon et al. 2021). Our method not only achieves state-of-the-art performance among deep learning approaches, but also surpasses well-known solvers and heuristic methods on ATSP instances by a significant margin.
Figure 1: The figure illustrates the CO problems we aim to address, and compares various learning-based approaches for CO problems. (Left) We focus on problems involving up to two distinct sets of items, which can be represented as a matrix. In ATSP, rows and columns represent cities (from/to), with each matrix element indicating distance. In PMSP, rows and columns represent machines and jobs, with each element representing processing time. (Center) The diagram highlights how these three approaches differ in generating solutions. Auto-regressive methods generate solutions by iterating through items, while diffusion-based create a heatmap followed by a problem-specific heatmap search process. Our proposed IC/DC approach generates feasible solutions directly, using a training procedure that guides the diffusion model to satisfy the constraints.
2 Related Works
For widely-studied CO problems like the travelling salesman problem (TSP), several off-the-shelf solvers are available, such as CPLEX (IBM 2022) and OR-tools (Google 2024). These solvers are build on a variety of heuristics, incorporating search methods (Helsgaun 2023), mathematical programming (Arora 1996), and graph algorithms (Christofides 2022). Typically, these approaches rely on problem-specific, handcrafted techniques, which limits their flexibility in adapting to diverse variants encountered in real-world scenarios.
Auto-regressive methods
To overcome this limitation, learning-based solvers have been developed, with early studies primarily focusing on auto-regressive approaches. Bello et al. (2016) were the first to propose solving CO problems using a pointer network trained via reinforcement learning (RL). Kwon et al. (2021) introduced an architecture called MatNet, which builds on the graph attention network (GAT) to encode various types of objects, enabling it to tackle more complex CO problems. While these auto-regressive models offer fast solution generation and can manage intricate CO problems, they are limited by their inability to revise previously made decisions and have been outperformed by methods utilizing diffusion-based methods.
Diffusion-based methods
To the best of our knowledge, two notable works have effectively addressed CO problems using diffusion models, both achieving solution quality comparable to off-the-shelf solvers while significantly reducing generation time. DIFUSCO (Sun and Yang 2023) is a graph neural network (GNN)-based diffusion model trained in a supervised manner to replicate solutions generated by solvers. Similarly, UTSP (Min, Bai, and Gomes 2024) employs a GNN-based diffusion model trained in an unsupervised manner, eliminating the need for costly solution generation from traditional solvers. However, UTSP’s objective is based on the concept of the Hamiltonian cycle, limiting its applicability to TSP.
However, these algorithms fall into the category of heatmap-generating approaches. While CO problems typically impose strict constraints on solutions, such as forming a Hamiltonian cycle, these algorithms are not trained to inherently satisfy those constraints. Instead, they rely on additional heatmap search techniques, such as active search methods (Qiu, Sun, and Yang 2022) or Monte Carlo Tree Search (MCTS) (Silver et al. 2016; Fu, Qiu, and Zha 2021), to produce feasible solutions. This reliance limits their applicability to CO problems with varying constraints on solutions.
3 Improving Combinatorial Optimization through Diffusion with Constraints
To combine the high-quality solutions of diffusion-based methods with the flexibility of auto-regressive approaches, we propose Improving Combinatorial optimization through Diffusion with Constraints (IC/DC). This approach ensures the feasibility of solutions while training diffusion model in a self-supervised manner, eliminating the need for costly supervision and problem-specific search processes.
3.1 Problem Definition
We consider a family of CO problems $\mathcal{C}$, which involve two distinct sets of items. Each problem $c\in\mathcal{C}$ is defined by two sets of items $\mathcal{A}$ and $\mathcal{B}$, and matrices that describe the relationships between these two sets of items, as illustrated on the left side of Figure 1. The solution to a CO problem $c$ is represented by a binary matrix $X\in\mathcal{X}=\{0,1\}^{|\mathcal{A}|\times|\mathcal{B}|}$.
Typically, for each problem $c$, there exists a feasible set of solutions, and a particular solution $X$ is evaluated using a problem-specific scoring function $\mathrm{score}:\mathcal{X}\times\mathcal{C}\rightarrow\mathbb{R}$ when it is feasible. For clarity, we define the reward function $R:\mathcal{X}\times\mathcal{C}\rightarrow\mathbb{R}$ as follows:
$R(X,c)=\begin{cases}\mathrm{score}(X,c)&\text{if $X$ is feasible for $c$}\\ -\infty&\text{otherwise}\end{cases},$
which allows us to express the objective of the CO problem $c$ as $\max_{X\in\mathcal{X}}\,R(X,c)$.
3.2 Training of IC/DC
We build upon a discrete diffusion model with categorical corruption processes (Austin et al. 2021). We represent the uncorrupted solution that we aim to generate as $X_{0}$, with the corrupted latent variables denoted as $X_{1},...,X_{T}$. We use lowercase $x$ to represent the vectorized forms of $X$, where $x_{t}=\operatorname{vec}(X_{t})\in\{0,1\}^{|\mathcal{A}||\mathcal{B}|}$, and tilded $\tilde{x}$ to denote the one-hot encoded versions, $\tilde{x}_{t}\in\{0,1\}^{|\mathcal{A}||\mathcal{B}|\times 2}$. In line with diffusion model conventions, we use $q(\cdot)$ to denote the data distribution/generative forward process, while $p_{\theta}(\cdot)$ represents the denoising reverse process, which is learned to generate the solutions.
Forward process
Our forward process is defined as:
$q(X_{t}|X_{t-1})$ $=\operatorname{Cat}(\operatorname{vec}^{-1}(\tilde{x}_{t-1}Q_{t})),$
$q(X_{t}|X_{0})$ $=\operatorname{Cat}(\operatorname{vec}^{-1}(\tilde{x}{Q}_{1:t})),$
$q(X_{t-1}|X_{t},X_{0})$ $=\operatorname{Cat}\left(\operatorname{vec}^{-1}\left(\frac{\tilde{x}_{t}Q_{t}^{\top}\odot\tilde{x}_{0}Q_{1:t-1}}{\tilde{x}_{0}Q_{1:t}\tilde{x}_{t}^{\top}}\right)\right),$
where $Q_{1:t}=Q_{1}Q_{2}...Q_{t}$, $\odot$ denotes the element-wise multiplication, and $\operatorname{vec}^{-1}$ reshapes the input to the shape $|\mathcal{A}|\times|\mathcal{B}|\times 2$. The matrix $Q_{t}\in[0,1]^{2\times 2}$ is a noise transition matrix that independently applies noise to each element of the solution matrix. We design this noise transition matrix to align with the prior distribution of feasible solutions $\bar{q}$ in the limit (Vignac et al. 2022), such that $\lim_{T\rightarrow\infty}Q_{1:T}z=\bar{q}$ for any vector $z$. This is achieved by defining:
$Q_{t}=\alpha_{t}I+\beta_{t}\mymathbb{1}\bar{q}^{\top},$
[…]
Reverse process
We follow the parametrization of Austin et al. (2021), where neural network $f_{\theta}(\cdot)$ is trained to directly predict logits of $X_{0}$ from each $X_{t}$, as follows:
$p_{\theta}(X_{0}|X_{t},c)$ $=\operatorname{Cat}(f_{\theta}(X_{t},t,c)),$
$p_{\theta}(X_{t-1}|X_{t},c)$ $\propto\sum_{X_{0}}q(X_{t-1},X_{t}|X_{0})p_{\theta}(X_{0}|X_{t},c).$
Although this parameterization facilitates the easy computation of diffusion loss when a target dataset is provided, samples from $p_{\theta}$ may not satisfy the constraints since each element of the solution matrix is independently sampled from a Categorical distribution. This limitation required the use of a feasibility-enforcing search process in previous diffusion-based studies on CO (Sun and Yang 2023; Min, Bai, and Gomes 2024). However, this approach requires a search process specifically tailored to each CO problem, and there is no assurance that the search process will preserve the quality of the solution that the diffusion model aims to generate.
Feasibility-enforced generation
To ensure the generation of feasible solutions, we design a process inspired by auto-regressive methods, which we call the feasibility-enforced generation process. In this approach, we sample one element of the solution matrix at a time, ensuring its feasibility based on the previously sampled elements, as follows:
$\widehat{p}_{\theta}(X_{0}|X_{t},c)$ $\propto\prod_{i=1}^{|\mathcal{A}||\mathcal{B}|}\operatorname{Cat}([X_{0}]_{i}|[f_{\theta}(X_{t},t,c)]_{i})\mathbb{I}([X_{0}]_{i}),$
$\mathbb{I}([X_{0}]_{i})$ =$=\begin{cases}1&\text{$[X_{0}]_{i}$ is feasible given $[X_{0}]_{1:i-1}$}\\ 0&\text{otherwise}\end{cases}.$
This approach, similar to the flexibility of auto-regressive methods, allows for the straightforward enforcement of feasibility in the generated samples. However, $\widehat{p}_{\theta}$ cannot be directly utilized as the reverse process because it involves discrete sampling, which prevents the use of the reparameterization trick, thereby making conventional and efficient variational training methods inapplicable. When using $p_{\theta}$ as the reverse process, the connection between $p_{\theta}$ and $\widehat{p}_{\theta}$ becomes weak, leading to a lack of guarantee that samples from $\widehat{p}_{\theta}$ will retain the desired characteristics.
Alternating training
To this end, we propose an iterative training approach that alternates between the CLONING step and the IMPROVEMENT step. In the CLONING step, we update the reverse process $p_{\theta}$ by maximizing the (lower bound of the) log likelihood of a set of high scoring feasible solutions: the surrogate targets. This guides $p_{\theta}$ toward generating high-quality feasible solutions, and strengthen its alignment with $\widehat{p}_{\theta}$, as they become identical when $p_{\theta}$ generates feasible samples only. In the IMPROVEMENT step, we directly update $\widehat{p}_{\theta}$ using reinforcement learning to maximize the scores of generated solutions. These two steps work in tandem, the IMPROVEMENT step is direct but computationally intensive, and CLONING step is more efficient but is only an indirect method of improving samples from $\widehat{p}_{\theta}$.
Surrogate targets
In standard diffusion model training, the target distribution $q(X_{0}|c)$, which represents the distribution of optimal solutions given a problem $c$, is typically available. However, in CO problems, obtaining such supervised dataset is often prohibitively expensive. To address this, we propose training our diffusion model in a self-supervised manner using a surrogate target distribution $\tilde{q}(X_{0}|c)$ instead.
This surrogate distribution is progressively refined during training and is defined as a reward-weighted mixture of two distributions:
$\tilde{q}(X_{0}|c)$ $\propto\exp(R(X_{0},c))[(1-\alpha)q(X_{0})+\alpha p_{\theta}(X_{0}|c)],$
where the mixture is controlled by the hyperparameter $\alpha\in[0,1]$. In the initial stage of training, the solutions generated by reverse process $p_{\theta}(X_{0}|c)$ are not feasible in general. This leads to using more samples from the prior distribution of feasible solutions $q(X_{0})$, and guide the diffusion model to more generate feasible solutions. As training progresses and as $p_{\theta}(X_{0}|c)$ begins to generate feasible solutions, the reward-weighting allows the diffusion model to refine itself by focusing on its high-scoring, feasible generations. Meanwhile, the inclusion of prior distribution of feasible solutions $q(X_{0})$ introduces diversity to the training, counteracting the tendency of $p_{\theta}$ to become too narrow as training progresses.
CLONING step
With the surrogate target distribution defined above, we perform standard diffusion training by minimizing the KL-divergence between the model’s generative distribution and the surrogate target distribution:
$\min_{\theta}D_{KL}(\tilde{q}(X_{0}|c)\parallel p_{\theta}(X_{0}|c)),$
which results in a variational bound objective (see Appendix A.2 for detailed derivations), $\mathcal{L}_{\text{VB}}(\theta):=$
$\mathbb{E}_{X_{0}\sim\tilde{q}}\left[\sum_{t=2}^{T}D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))\right]$
Inspired by recent practices (Austin et al. 2021), we also incorporate the following auxiliary losses:
$\mathcal{L}_{\text{prd}}(\theta)$ $:=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[-\log p_{\theta}(X_{0}|X_{t},c)\right],$
$\mathcal{L}_{\text{cst}}(\theta)$ $:=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[C(p_{\theta}(X_{0}|X_{t},c))\right],$
where $\mathcal{L}_{\text{prd}}$ encourages accurate predictions of the data $X_{0}$ at each time step, and $\mathcal{L}_{\text{cst}}$ discourages infeasible predictions, with $C(\cdot)$ being a differentiable function that approximately measures constraint violations of samples using the Gumbel-softmax trick (see Appendix. A.3 for details on $C$).
In summary, during the CLONING step, we minimize:
$\mathcal{L}_{\texttt{CLN}}(\theta)=\mathcal{L}_{\text{VB}}(\theta)+\lambda_{1}\mathcal{L}_{\text{prd}}(\theta)+\lambda_{2}\mathcal{L}_{\text{cst}}(\theta).$ (1)
IMPROVEMENT step
To directly improve the feasibility-enforced generations, we minimize the following objective:
$\mathcal{L}_{\texttt{IMP}}(\theta)=-\mathbb{E}_{X_{0}\sim\widehat{p}_{\theta}(X_{0}|c)}[R(X_{0},c)].$
Similar to auto-regressive methods, the feasibility-enforced generation process can be viewed as a sequential decision making task, where each element of the solution matrix $X_{0}$ is determined step by step. This perspective allows us to compute the gradient of the above objective using the REINFORCE algorithm (Williams 1992).
After sampling a set of solutions $\{X_{0}^{(1)},X_{0}^{(2)},...,X_{0}^{(N)}\}$ using $\widehat{p}_{\theta}(X_{0}|c)$, we approximate the gradient as follows: $\nabla_{\theta}\mathcal{L}_{\texttt{IMP}}(\theta)\approx$
$-\frac{1}{N}\sum_{i=1}^{N}\left(R^{(i)}-\frac{1}{N}\sum_{j=1}^{N}R^{(j)}\right)\nabla_{\theta}\log\widehat{p}_{\theta}\left(X_{0}^{(i)}|c\right),$ (2)
where $R^{(i)}=R(X_{0}^{(i)},c)$. This approach is adapted from the baseline estimation method of POMO (Kwon et al. 2020).
Summary
We train IC/DC by alternating between the CLONING step–diffusion model training with surrogate targets–and the IMPROVEMENT step–reinforcement learning of feasibility-enforced generation, as shown in Algorithm 1. In practice, we use a replay memory $\mathcal{D}_{\tilde{q}}$ to implement a surrogate target distribution, and perform multiple CLONING steps for each IMPROVEMENT step. This is because CLONING updates only a single timestep of the diffusion model, and IMPROVEMENT is significantly slower due to the online generations.
Algorithm 1 Training IC/DC
Input: A set of CO problems $\mathcal{C}$, learning late $\gamma$, diffusion step $T$, target mix ratio $\alpha$
1:  function CLONING(${\theta}$)
2:     $X_{0},c\sim\mathcal{D}_{\tilde{q}}(\alpha)$ with a probability proportional to $\exp(R(X_{0},c))$
3:     $t\sim\text{Uniform}(\{1,...,T\})$
4:     $X_{t}\sim q(X_{t}|X_{0})$
5:     $\theta\leftarrow\theta-\gamma\nabla_{\theta}\mathcal{L}_{\texttt{CLN}}$ according to Eq. (1)
6:  end function
6:
7:  function IMPROVEMENT(${\theta}$)
8:     for $i\leftarrow 1$ to $N$ do
9:        Sample $c^{(i)}$ from a set of CO problems $\mathcal{C}$
10:        $X_{T}^{(i)}\sim q(X_{T}^{(i)})$
11:        for $t\leftarrow T$ to $1$ do
12:           ${X}_{t-1}^{(i)}\sim p_{\theta}(X_{t-1}|{X}_{t}^{(i)},c^{(i)})$
13:        end for
14:        $\widehat{X}_{0}^{(i)}\sim\widehat{p}_{\theta}(X_{0}|X_{1}^{(i)},c^{(i)})$
15:     end for
16:     $\theta\leftarrow\theta-\gamma\nabla_{\theta}\mathcal{L}_{\texttt{IMP}}(\theta)$ with $\{(\widehat{X}_{0}^{(i)},c^{(i)})\}_{i=1}^{N}$ according to Eq. (2)
17:     Sample $\{(\tilde{X}_{0}^{(i)},\tilde{c}^{(i)})\}_{i=1}^{\frac{1-\alpha}{\alpha}N}$ from $q(X_{0})$, $\mathcal{C}$
18:     Store $\{(X_{0}^{(i)},c^{(i)})\}_{i=1}^{N},\{(\tilde{X}_{0}^{(i)},\tilde{c}^{(i)})\}_{i=1}^{\frac{1-\alpha}{\alpha}N}$ in $\mathcal{D}_{\tilde{q}}$
19:  end function
19:
20:  Initialize an empty set $\mathcal{D}_{\tilde{q}}=\{\}$
21:  repeat
22:     for  $m\leftarrow 1$ to $M$ do
23:        CLONING(${\theta}$, $\mathcal{D}_{\tilde{q}}$)
24:     end for
25:     IMPROVEMENT(${\theta}$, $\mathcal{D}_{\tilde{q}}$)
26:  until convergence
Figure 2: This figure illustrates our architecture and diffusion process. (Bottom-right) The problem encoder encodes a problem instance $c=(A,B,D)$ into a problem embedding $(A^{\prime},B^{\prime})$. (Bottom-left) The denoiser takes a problem embedding, noisy solution and a timestep embedding $(A^{\prime},B^{\prime},X_{t},t)$ and outputs denoised solution $X_{0}$.
4 Architecture
The neural network we need is $f:\mathcal{X}\times\mathcal{C}\rightarrow\mathcal{X}$, which encodes the CO problem and outputs a distribution over binary matrices given an input binary matrix. To achieve this, we propose a problem encoder that effectively encodes CO problem, and a denoiser, a specialized variant of GNN that processes the bipartite graph between two sets of items.
Problem encoder
For the CO problems we consider, a problem instance $c$ consists of information about two sets of items and their relationships. For simplicity, let’s assume that all items share the same number of features $d$; if not, different embedding layers can be used to standardize the feature dimensions. The information for the items in set $\mathcal{A}$ is represented by the matrix $A\in\mathbb{R}^{|\mathcal{A}|\times d}$, and similarly, the items in set $\mathcal{B}$ are represented by the matrix $B\in\mathbb{R}^{|\mathcal{B}|\times d}$. There relationship between these items is captured by the matrix $D\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|}$. Together, these matrices define the problem instance, i.e., $c=(A,B,D)$.
To effectively encode the problem represented by these matrices, we adopt the dual graph attentional layer structure from MatNet (Kwon et al. 2021), but replace the attention layer with a modified version of graph attention networks (GAT, Veličković et al. 2017) that is specifically designed to process a bipartite graph. The problem encoder consists of $L$ layers, where each layer takes $(A,B,D)$ as input and outputs updated features $(A^{\prime},B^{\prime})$. The outputs of the final layer are then passed to the denoiser. The bottom-right side of Figure 2 illustrates this problem encoder. Detailed equations of problem encoder layer are provided in Appendix B.2.
Denoiser
Building on recent empirical successes (Joshi et al. 2020; Qiu, Sun, and Yang 2022), we extend the anisotropic graph neural network (AGNN) to handle bipartite graphs, allowing us to consider two distinct sets of items, and use it as the denoiser. The input embedding layer maps each element of the noisy solution matrix $X_{t}$ and the timestep $t$ into $d$-dimensional features. These embeddings are then passed to the AGNN, along with the problem embedding $A^{\prime}$ and $B^{\prime}$. After $L^{\prime}$ layers of AGNN, the embedded solution matrix with updated features is passed through a linear layer to produce the denoised solution matrix $X_{0}$. The bottom-left side of Figure 2 illustrates this denoiser. Detailed equations of denoiser layer are provided in Appendix B.3.
5 Experiments
PMSP-20 PMSP-50
Makespan $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Makespan $\downarrow$ Gap $\downarrow$ Time $\downarrow$
Baseline CP-SAT (OR-tool) 28.110 0% (1m) 65.900 0% (3m)
(Meta-)heuristics Random 59.468 71.612% (1m) 138.683 71.153% (3m)
Shortest Job First 34.250 19.692% (1m) 72.986 10.204% (2m)
Genetic Algorithm 30.211 7.205% (2h) 68.912 4.468% (4h)
Particle Swarm Opt. 29.794 5.817% (1h) 69.133 4.788% (3h)
Auto-regressive MatNet 29.588 5.123% (0s) 67.092 1.793% (1s)
MatNet (×$\times 8$) 28.491 1.346% (1s) 66.129 0.347% (3s)
MatNet (×$\times 32$) 28.353 0.860% (2s) 66.105 0.311% (16s)
MatNet (×$\times 128$) 28.283 0.615% (15s) 66.020 0.182% (1m)
[…]
Ours IC/DC 29.168 3.694% (7s) 66.677 1.172% (10s)
IC/DC (×$\times 8$) 28.300 0.674% (15s) 66.196 0.448% (28s)
IC/DC (×$\times 32$) 28.190 0.284% (46s) 66.055 0.235% (1m)
IC/DC (×$\times 128$) 28.150 0.142% (3m) 65.974 0.112% (5m)
Table 1: (Results for PMSP-20 and PMSP-50) The top row (labeled as baseline) represents the results from an off-the-shelf solver and is used as a reference for calculating the performance gap. $\infty$ indicates that the generated solutions are infeasible. Results where our model outperforms the previous SOTA learning-based method are highlighted in bold.
ATSP-20 ATSP-50
Tour Length $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Tour Length $\downarrow$ Gap $\downarrow$ Time $\downarrow$
Baseline CPLEX 1.534 0% (2m) 1.551 0% (2h)
(Meta-)heuristics Nearest Neighbor 1.994 26.099% (0s) 2.092 29.701% (0s)
Nearest Insertion 1.791 15.477% (0s) 1.938 22.164% (0s)
Furthest Insertion 1.709 10.770% (0s) 1.836 16.802% (0s)
LKH-3 1.561 1.758% (1s) 1.551 0% (8s)
Auto-regressive MatNet 1.541 0.456% (0s) 1.572 1.314% (1s)
MatNet (×$\times 8$) 1.535 0.078% (2s) 1.556 0.317% (15s)
MatNet (×$\times 32$) 1.534 0.033% (9s) 1.554 0.214% (31s)
MatNet (×$\times 128$) 1.534 0.013% (37s) 1.553 0.111% (2m)
[…]
Ours IC/DC 1.609 4.802% (1s) 1.614 3.969% (15s)
IC/DC (×$\times 8$) 1.549 0.971% (8s) 1.566 0.928% (2m)
IC/DC (×$\times 32$) 1.534 -0.026% (31s) 1.547 -0.235% (8m)
IC/DC (×$\times 128$) 1.530 -0.235% (2m) 1.543 -0.532% (30m)
Table 2: (Results for ATSP-20 and ATSP-50) The top row (labeled as baseline) represents the results from an off-the-shelf solver and is used as a reference for calculating the performance gap. $\infty$ indicates that the generated solutions are infeasible.
5.1 Demonstrated Problems
We begin by describing the combinatorial optimization (CO) problems on which we conducted experiments: the Parallel Machine Scheduling Problem (PMSP) and the Asymmetric Travelling Salesman Problem (ATSP).
Parallel machine scheduling problem
In PMSP, a problem instance $c$ consists of $|\mathcal{J}|$ jobs and $|\mathcal{M}|$ machines. Each job $j\in\mathcal{J}$ must be scheduled on a machine $m\in\mathcal{M}$, with varying workloads for each job and different processing capabilities for each machine. The primary objective in PMSP is to minimize the makespan, which is the total length of the schedule upon the completion of all jobs. In this context, having $[X_{0}]_{j,m}=1$ indicates that job $j$ is assigned to machine $m$, which takes a processing time of $[P]_{j,m}$ where $P\in\mathbb{R}_{+}^{|\mathcal{J}|\times|\mathcal{M}|}$ is a matrix of processing times for all combinations. The goal is to determine the solution matrix $X_{0}=\{0,1\}^{|\mathcal{J}|\times|\mathcal{M}|}$ that minimizes the makespan for a given problem $c=(\mathcal{M},\mathcal{J},P)$:
$\text{score}_{\text{PMSP}}(X_{0},c)=-\max_{m}\sum_{j}[X_{0}\odot P]_{j,m}.$
The solution matrix that assigns a job to multiple machines is considered infeasible.
Asymmetric travelling salesman problem
An ATSP instance $c=(|\mathcal{N}|,D)$ comprises $|\mathcal{N}|$ cities and an asymmetric distance matrix $D\in\mathbb{R}_{+}^{|\mathcal{N}|\times|\mathcal{N}|}$ where each element of it specifies the distance between two cities. The solution to ATSP is a tour, which is an adjacency matrix $X_{0}=\{0,1\}^{|\mathcal{N}|\times|\mathcal{N}|}$ for a directed graph visiting all cities once. The goal is find a solution that minimizes the tour length:
$\text{score}_{\text{ATSP}}(X_{0},c)=-\sum_{i,j}{[X_{0}\odot D]_{i,j}}.$
The solution matrix that is not a Hamiltonian cycle is considered infeasible. In accordance with Kwon et al. (2021) we employ tmat-class ATSP instances (see Appendix D.1).
5.2 Main Results
We evaluated the baselines and the proposed algorithm using two Intel Xeon Gold 6330 CPUs and an RTX 3090 GPU for both PMSP and ATSP. For the evaluation, 1000 problem instances were randomly generated using a standard generation process (see appendix D.1). To fully leverage the stochastic nature of generative learning-based methods, we also evaluated these methods by generating multiple samples (×$\times n$) for each problem instance and selecting the one with the best score. For MatNet (Kwon et al. 2021), we followed the authors’ implementation, including instance augmentation, which yielded better result. As a problem-specific search process has not been studied on PMSP and ATSP, for diffusion-based methods, we report simple discrete diffusion models, either trained with supervised learning (Sun and Yang 2023) or with reinforcement learning (Black et al. 2023). For further experimental details, please refer to Appendix C and Appendix D.
PMSP
We evaluated our method on PMSP-20 and PMSP-50, where the numbers 20 and 50 correspond to the number of jobs in each instance, with the number of machines fixed at 4. Our approach is compared against various baselines, including (meta-)heuristics, auto-regressive, and diffusion-based methods. Details of these baselines can be found in Appendix C.2.
As shown in Table 1, IC/DC achieves the smallest optimality gap among learning-based methods. IC/DC demonstrates a $0.142\%$ performance gap compared to CP-SAT, whereas the previous SOTA, MatNet, shows a $0.615\%$ gap on PMSP-20. On PSMP-50 IC/DC reduces the gap from MaNet’s $0.182\%$ to $0.112\%$. Although IC/DC has a slower inference speed compared to auto-regressive methods, its high-quality solutions remain highly competitive with other baselines.
ATSP
We evaluated our method on ATSP-20 and ATSP-50, where the number 20 and 50 correspond to the number of cities. IC/DC generates solutions that not only surpass those generated by SOTA learning-based methods but also outperform those obtained through (meta-)heuristics and off-the-shelf solvers. IC/DC achieves a $-0.235\%$ gap on ATSP-20 and a $-0.532\%$ gap on ATSP-50, demonstrating a negative gap compared to a powerful CPLEX solver—which no other learning-based method has achieved. This substantial improvement positions IC/DC as a new SOTA method for addressing these less-studied CO problems, highlighting its potential.
5.3 Discussion
Comparison with auto-regressive method
As discussed earlier, the performance of auto-regressive models is limited by their sequential decision-making process, which locks in previously made decisions, preventing any revisions. While IC/DC employs a similar auto-regressive approach for feasibility-enforced generation, it also benefits from the accuracy of diffusion-based methods by iteratively refining the solution through $T-1$ denoising steps before the final auto-regressive generation. Moreover, this two-stage generation process enables IC/DC to achieve much greater sample diversity compared to auto-regressive models, as the denoising process provides varied foundations for solutions before the auto-regressive generation.
The results that support these arguments can be found in both tables, particularly in Table 2. While IC/DC’s performance lags behind MatNet when using a single generation to solve the problem, its quickly surpasses MatNet as the number of samples increases, highlighting the high diversity of IC/DC’s samples. Upon closer examination, we observed that the stochasticity of MatNet primarily stems from the starting point of the auto-regressive generation (i.e., the selection of the initial city to begin the tour). In contrast, IC/DC’s denoising process offers diverse backbones for the auto-regressive generation, resulting in varied samples even when starting from the same initial city.
Comparison with diffusion-based method
Previously proposed diffusion-based methods (Qiu, Sun, and Yang 2022; Sun and Yang 2023; Min, Bai, and Gomes 2024) have employed problem-specific search processes to enforce feasibility and enhance performance. However, developing such search processes with strong performance and feasibility guarantees typically requires significant expertise in the specific CO problem being addressed. As the complexity of CO problem increases, such as with ATSP or more complex TSP variants compared to TSP, the cost of implementing these search algorithms becomes even greater.
As shown in Table 1 and Table 2, simple implementations of diffusion models, whether trained with supervised learning (Sun and Yang 2023) or reinforcement learning (Black et al. 2023), struggle to generate feasible solutions without the aid of problem-specific search processes, which have not been studied for these particular CO problems. This highlights the critical importance of our feasibility-enforced generation process, which integrates the flexibility of auto-regressive methods into diffusion-based methods, allowing for the straightforward imposition of various constraints. In this regard, the proposed IC/DC approach not only demonstrates superior performance compared to previous learning-based methods, but also significantly broadens the applicability of diffusion-based methods to a wide range of less-explored CO problems with diverse constraints.
6 Conclusion
We propose a diffusion-based approach named IC/DC to tackle the CO problems involving two distinct sets of items. Our algorithm demonstrates superior performance in both PMSP and ATSP, outperforming existing baselines and successfully generating feasible solutions that cannot be easily achieved with simple diffusion-based methods. We emphasize that this is the first study to successfully adapt a diffusion-based method for CO problems involving two distinct sets of items. We believe the proposed IC/DC framework has significant potential for generlizability. For example, in Appendix D.4, we demonstrate its capability to address real-world CO problems with sophisticated item and relationship features.
Despite its strong performance, IC/DC has a notable limitation: the GAT-based encoder we use requires $O(\max(|\mathcal{A}|,|\mathcal{B}|)^{2})$ memory complexity. This demands substantial memory resources when training on large instances. To address these challenges, we plan to explore memory-efficient techniques in future work, such as those introduced by Zhu et al. (2024).
Appendix A Derivation
A.1 Joint Variational Upper Bound with EUBO
We derive the upper bound using the Evidence Upper Bound (EUBO) (Ji and Shen 2019)
[…]
The inequality is established by the Gibbs’ inequality: $-\int\tilde{q}(Z|X,c)\log\tilde{q}(Z|X,c)\leq-\int\tilde{q}(Z|X,c)\log p_{\theta}(Z|X,c)$.
$D_{KL}(\tilde{q}(X|c)\|p_{\theta}(X|c))$ $=\int\tilde{q}(X|c)\log\frac{\tilde{q}(X|c)}{p_{\theta}(X|c)}dX$
[…]
$=D_{KL}(\tilde{q}(X,Z|c)||p_{\theta}(X,Z|c))$
[…]
A.2 KL divergence to Reward-weighted Diffusion Loss
We derive the reward-weighted diffusion loss from the objective in CLONING step. We aim to minimize the Kullback-Leibler (KL) divergence $D_{KL}(\tilde{q}(X|c)\parallel p_{\theta}(X|c))$ between the surrogate target and the distribution $p$ parameterized by $\theta$. however since evaluating the log-likelihood of the generative model $p_{\theta}(X|c)$ is difficult (Kingma and Welling 2013), we instead utilize a joint variational upper bound (see App. A.1):
$\mathrm{D}_{\mathrm{KL}}(\tilde{q}(X|c)\parallel p_{\theta}(X|c))\leq\mathrm{D}_{\mathrm{KL}}(\tilde{q}(X,Z|c)\parallel p_{\theta}(X,Z|c))$
According to Gibbs’ inequality if the right-hand side of the inequality is zero, the inequality becomes an equality. In this case, the surrogate target is exactly approximated.
Therefore, we focus on minimizing the right-hand side of eq (A.2). By applying the diffusion process where $Z=X_{1:T}$, $X=X_{0}$, and $T$ is the diffusion step, the forward process is defined by $q(X_{1:T}|X_{0})$ with $X_{0}\sim\tilde{q}(X_{0}|c)$. We minimize the Reward-weighted Diffusion Loss $\mathcal{L}_{\text{RWD}}$:
$\mathcal{L}_{\text{RWD}}$ $=D_{KL}(\tilde{q}(X_{0:T}|c)\parallel p_{\theta}(X_{0:T}|c))$
$=D_{KL}(\tilde{q}(X_{0},X_{1:T}|c)\parallel p_{\theta}(X_{0:T}|c))$
$=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{1:T}\sim q(X_{1:T}|X_{0})}\left[\log\frac{q(X_{1:T}|X_{0})\tilde{q}(X_{0}|c)}{p_{\theta}(X_{0:T}|c)}\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{q(X_{1:T}|X_{0})}{p_{\theta}(X_{0:T}|c)}+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{\prod_{t=1}^{T}q(X_{t}|X_{t-1})}{p_{\theta}(X_{T}|c)\prod_{t=1}^{T}p_{\theta}(X_{t-1}|X_{t},c)}+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[-\log p_{\theta}(X_{T}|c)+\sum_{t=1}^{T}\log\left(\frac{q(X_{t}|X_{t-1})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[-\log p_{\theta}(X_{T}|c)+\sum_{t=2}^{T}\log\left(\frac{q(X_{t}|X_{t-1})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)+\log\left(\frac{q(X_{1}|X_{0})}{p_{\theta}(X_{0}|X_{1},c)}\right)+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\Bigg{[}-\log p_{\theta}(X_{T}|c)$
$\quad\quad\quad\quad\quad\quad\quad\ \ +\sum_{t=2}^{T}\log\left(\frac{q(X_{t-1}|X_{t},X_{0})}{p_{\theta}(X_{t-1}|X_{t},c)}\cdot\frac{q(X_{t}|X_{0})}{q(X_{t-1}|X_{0})}\right)+\log\left(\frac{q(X_{1}|X_{0})}{p_{\theta}(X_{0}|X_{1},c)}\right)+\log\tilde{q}(X_{0}|c)\Bigg{]}$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{q(X_{T}|X_{0})}{p_{\theta}(X_{T}|c)}+\sum_{t=2}^{T}\log\left(\frac{q(X_{t-1}|X_{t},X_{0})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)-\log p_{\theta}(X_{0}|X_{1},c)+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\Bigg{[}\underbrace{D_{KL}(q(X_{T}|X_{0})\|p_{\theta}(X_{T}|c))}_{L_{T}}+\sum_{t=2}^{T}\underbrace{D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))}_{L_{t-1}}$
$\quad\quad\quad\quad\quad\quad\quad\ \ +\underbrace{D_{KL}(\tilde{q}(X_{0}|c)\|p_{\theta}(X_{0}|X_{1},c))}_{L_{0}}\Bigg{]}$
If $T$ is sufficiently large, $L_{T}$ will approach zero (Austin et al. 2021). Also, $L_{0}$ can be derived as CE loss:
$L_{0}$ $=\sum\tilde{q}(X_{0}|c)\log\frac{\tilde{q}(X_{0}|c)}{p_{\theta}(X_{0}|X_{1},c)}=\sum\tilde{q}(X_{0}|c)\log\tilde{q}(X_{0}|c)-\log p_{\theta}(X_{0}|X_{1},c))$
$=CE(\tilde{q},p_{\theta})-H(\tilde{q})\approx\mathcal{L}_{\text{prd}}-H(\tilde{q})$
Cross-entropy loss is equal to $\mathcal{L}_{\text{prd}}$ at $t=1$. As $H(\tilde{p})$ is independent to $p_{\theta}$, it can be ignored for training. As $f_{\theta}$ predicts the logit of $X_{0}$, we use the Cross-entropy loss for every timestep $L_{\text{pred}}$ instead of Cross-entropy loss at $t=1$.
$\therefore\mathcal{L}_{\text{RWD}}\approx\mathcal{L}_{\text{VB}}+\mathcal{L}_{\text{prd}}$
Training with this loss function results in an enhancement of solution quality.
A.3 Constraint Loss
In CO problems, solutions must strictly satisfy specific constraints. To discourages infeasible predictions $X_{0}\sim p_{\theta}(X_{0}|X_{t},c)$, we introduce a constraint loss:
$\mathcal{L}_{\text{cst}}(\theta)$ $=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[C(p_{\theta}(X_{0}|X_{t},c))\right],$
where $C(\cdot)$ is a differentiable function that approximately measures constraint violations of samples
In the Asymmetric Travelling Salesman Problem (ATSP), each node is required to travel to a distinct city other than itself. Consequently, in the resulting solution matrix, each row and each column must contain exactly one entry of ’1’:
$C_{\text{ATSP}}(X_{0})=\left(\sum_{i}\text{Gumbel-Softmax}(X_{0})_{i,j,1}-1\right)^{2}+\left(\sum_{j}\text{Gumbel-Softmax}(X_{0})_{i,j,1}-1\right)^{2}$ (3)
In the Parallel Machine Scheduling Problem (PMSP), each job is required to be assigned to a single machine. Accordingly, in the solution matrix, each column must contain exactly one entry of ’1’, with all other entries in that column being ’0’:
$C_{\text{PMSP}}(X_{0})=\left(\sum_{j}\text{Gumbel-Softmax}(X_{0})_{i,j,1}-1\right)^{2}$ (4)
[…]
A.4 Diffusion Loss
By combining reward-weighted diffusion loss A.2 and constraint loss A.3, our diffusion objective:
$\mathcal{L}_{\texttt{CLN}}(\theta)=\mathbb{E}_{\tilde{q}}\left[\left(\sum_{t=2}^{T}D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))-\lambda_{1}p_{\theta}(X_{0}|X_{t},c)\right)+\lambda_{2}C(p_{\theta}(X_{0}|X_{t},c))\right]$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are hyper parameter.
Appendix B IC/DC in detail
B.1 Choice of Transition Matrices
According to (Austin et al. 2021), it is argued that incorporating domain-specific structures into the transition matrices $Q_{t}$ within the diffusion process is a reasonable approach. In our case, due to the inherent sparsity in the solution matrices of CO problems, the marginal distribution of feasible solutions significantly deviates from the uniform distribution commonly used in standard diffusion processes. Therefore, we design this noise transition matrix to align with the prior distribution of feasible solutions $\bar{q}$.
Depending on what CO problem we are aiming to solve, we are often able to compute the prior distribution, averaged over solution elements $\bar{q}(\tilde{x})=\bar{q}\left(\frac{1}{|\mathcal{A}||\mathcal{B}|}\sum_{i=1}^{|\mathcal{A}||\mathcal{B}|}[\tilde{x}_{0}]_{i}\right)$. The marginal probability of $x=1$ over a set of feasible solutions can be expressed as $\bar{q}(x=1)$.
In ATSP, when considering $|\mathcal{N}|$ cities, the solution involves travelling through all $|\mathcal{N}|$ cities. The prior distribution follows the following:
$\bar{q}=[1-\frac{1}{|\mathcal{N}|},\frac{1}{|\mathcal{N}|}]^{\top}$
In PMSP, when considering $|\mathcal{J}|$ jobs and $|\mathcal{M}|$ machines, the solution involves assigning all $|\mathcal{J}|$ jobs to the machines. The prior distribution is as follows:
$\bar{q}=[1-\frac{1}{|\mathcal{J}|},\frac{1}{|\mathcal{J}|}]^{\top}$
In most combinatorial optimization problems, the marginal distribution $\bar{q}$ is typically known. However, in cases where the marginal distribution is not available, an alternative approach is to utilize a uniform distribution. As a substitute for $\bar{q}$, one may consider using $\bar{u}=[0.5,0.5]$.
B.2 Problem Encoder
Denoting each row of $A$ as $A=[a_{1},...,a_{|\mathcal{A}|}]^{\top}$, an attention block within each layer processes the input as:
$S_{\text{inter}}$ $=\text{softmax}(AW_{\text{inter}}A^{\top})$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{A}|},$
$S_{\text{intra}}$ $=\text{ReLU}(AW_{\text{intra}}{B}^{\top})$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|},$
$S$ $=\text{concat}(S_{\text{inter}}S_{\text{intra}},D)$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|\times 2},$
$\tilde{D}$ $=\text{MLP}(S)$ $\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{B}|},$
$\tilde{A}$ $=\text{softmax}(\tilde{D})BW_{\text{v}}$ $\in\mathbb{R}^{|\mathcal{A}|\times d},$
where $W_{\text{inter}}$, $W_{\text{intra}}$, and $W_{\text{v}}$ are weight matrices of dimension $\mathbb{R}^{d\times d}$, and $\text{MLP}(\cdot)$ is a fully-connected neural network that maps $2$-dimensional inputs to $1$-dimensional outputs. The matrices $S_{\text{inter}}$ and $S_{\text{intra}}$ are designed to capture the inter-relationships within set $\mathcal{A}$ and the intra-relationships and between sets $\mathcal{A}$ and $\mathcal{B}$. These matrices, along with $D$, are combined to form the attention score $\tilde{D}$, which produces the output $\tilde{A}$. Using the described attention block, the layer outputs $A^{\prime}$ as follows:
$\hat{A}=\text{BN}(A+\tilde{A}),\quad A^{\prime}=\text{BN}(\hat{A}+\text{MLP}(\hat{A})),$
where BN refers to a batch normalization (Ioffe and Szegedy 2015). The process for updating $B$ to $B^{\prime}$ is computed in the same way.
B.3 Denoiser
As illustrated in the bottom-left side of Figure 2, the denoiser consists of $L^{\prime}$ layers, where each layer gets input of $(A,B,X,t)$ and outputs $(A^{\prime},B^{\prime},X^{\prime})$, which is processed as follows:
$\hat{x}_{i,j}^{\ell+1}$ $=P^{\ell}x_{i,j}^{\ell}+Q^{\ell}h_{i}^{\ell}+R^{\ell}h_{j}^{\ell},$
$x_{i,j}^{\ell+1}$ $=x_{i,j}^{\ell}+\text{MLP}_{x}(\text{BN}(\hat{x}_{i,j}^{\ell+1}))+\text{MLP}_{t}(\mathbb{t}),$
$h_{i}^{\ell+1}$ $=h_{i}^{\ell}+\text{ReLU}\left(\text{BN}(U^{\ell}_{a}h_{i}^{\ell}+\sum_{j\in\mathcal{N}_{i}}(\sigma(\hat{x}_{i,j}^{\ell+1})\odot V^{\ell}_{b}h_{j}^{\ell}))\right),$
$h_{j}^{\ell+1}$ $=h_{j}^{\ell}+\text{ReLU}\left(\text{BN}(U^{\ell}_{b}h_{j}^{\ell}+\sum_{i\in\mathcal{N}_{j}}(\sigma(\hat{x}_{j,i}^{\ell+1})\odot V^{\ell}_{a}h_{i}^{\ell}))\right),$
Where $h^{\ell=0}_{i}=h^{L}_{i}$ and $h^{\ell=0}_{j}=h^{L}_{j}$. For simplicity, the vector representation $x^{\ell=0}_{i,j}$ at the $(i,j)$-th $X_{t}$ is denoted without $t$.
The matrices $U^{\ell}_{a},U^{\ell}_{b},V^{\ell}_{a},V^{\ell}_{b},P^{\ell},Q^{\ell},R^{\ell}\in\mathbb{R}^{d\times d}$ are learnable parameters of the $\ell$-th layer. SUM pooling is denoted by $\sum$ (Xu et al. 2018), the sigmoid function is represented by $\sigma$, and the Hadamard product is denoted by $\odot$. $\mathcal{N}_{i}$ denotes the neighborhood of node $i$ among the $B$ items, while $\mathcal{N}_{j}$ denotes the neighborhood of node $j$ among the $A$ items. Additionally, the variable $\mathbb{t}$ denotes the sinusoidal features (Vaswani et al. 2017) corresponding to the denoising timestep $t$. After the final layer $\mathfrak{L}$, the output $x^{\mathfrak{L}}_{i,j}$ is passed through a linear layer to obtain clean matrix $X_{0}=\{x_{0,0},...,x_{I,J}\}$.
Appendix C PMSP definition and baselines
C.1 Unrelated Parallel Machine
In the literature on the parallel machine scheduling problem, most studies focus on a simplified variant where the processing time for a job is consistent across all machines, commonly referred to as ”uniform parallel machines”. In contrast, we examine a more complex scenario where the processing times for each machine are entirely independent to one another. For our study, we generate the processing time matrix randomly and use it as our instance.
C.2 Baseline
Mixed-integer programming
Constraint Programming with Satisfiability (CP-SAT) is a highly efficient solver developed as part of the OR-Tools suite, designed for solving integer programming problems. It is particularly effective for solving scheduling problems, including PMSP. For PMSP our MIP model is based on Avalos-Rosales, Alvarez, and Angel-Bello (2013)
$\text{minimize}\quad C_{max}$ (6)
s.t. $C_{0}=0$ (7)
$\sum_{j\in J}x_{i0j}\leq 1\quad\quad i\in M$ (8)
$\sum_{j\in J_{0},j\neq k}\sum_{i\in M}x_{ijk}=1\quad\quad k\in J$ (9)
$\sum_{k\in J_{0},j\neq k}\sum_{i\in M}x_{ijk}=1\quad\quad j\in J$ (10)
$\sum_{j\in J_{0},j\neq k}\sum_{k\in J}p_{ik}x_{ijk}\leq C_{max}\quad i\in M$ (11)
$\sum_{k\in J_{0},j\neq k}x_{ijk}=\sum_{h\in J_{0},h\neq j}x_{ihj}\quad\quad j\in J,i\in M$ (12)
$C_{k}-C_{j}+V(1-x_{ijk})\geq p_{ik}\quad j\in J_{0},k\in J,j\neq k,i\in M$ (13)
where
[…]
$C_{j}:\text{Completion time of job j}$
$x_{ijk}:\begin{cases}1&\text{if job $k$ is processed directly after job $j$ on machine $i$}\\ 0&\text{otherwise}\end{cases}.$
$p_{ik}:\text{Time required to process job $j$ on machine $i$}$
$M:\text{Set of machines}$
$J:\text{Set of jobs to schedule}$
[…]
$V:\text{A large positive number}$
Constraints (9) and (10) ensure that each job has exactly one predecessor on one of the machines. Constraint(12) specifies that if a job has a predecessor on a machine, it must also have a successor on that same machine. Constraint (13) guarantees that a valid sequence of jobs is scheduled on each machine, with no overlap in processing times. Constraint (8) ensures that only one job can be scheduled as the first job on each machine. Constraint (7) sets the completion time of job 0, an auxiliary job used to define the start of the schedule, to zero. Finally, constraint (11) establishes the relationship between the makespan of individual machines and the overall schedule makespan. CP-SAT is run on CPUs.
(Meta-)Heuristics
Random and Shortest Job First (SJF) are greedy-selection algorithms designed to generate valid schedules using the Gantt chart completion strategy. SJF specifically prioritizes tasks by scheduling the shortest available jobs in ascending order at each time step $t$. Although simple, these methods can serve as effective baselines, providing a comparison point for evaluating more sophisticated scheduling algorithms. The greedy-selection algorithms run on CPUs.
CO problems often require sophisticated methods to find high-quality solutions, particularly when traditional approaches like Mixed Integer Programming (MIP) are impractical due to complexity. In such cases, meta-heuristics provide a robust alternative. Genetic Algorithm (GA) and Particle Swarm Optimization (PSO) are two widely adopted meta-heuristics, known for their versatility and effectiveness across a range of problem domains.
GA iteratively updates multiple candidate solutions, referred to as chromosomes. New child chromosomes are produced by combining two parent chromosomes through crossover methods, and mutations are applied to the chromosomes to enhance exploration.
PSO iteratively updates multiple candidate solutions, referred to as particles. For every iteration each particles are updated based on the local best known and the global best known particles.
We utilize the implementations provided by Kwon et al. (Kwon et al. 2021) and follow their setting. Both GA and PSO run on GPUs.
MatNet
MatNet, proposed by (Kwon et al. 2021), adapts the attention model (Kool, Van Hoof, and Welling 2018) to be applicable to bipartite graphs. It employs cross-attention mechanisms in place of self-attention to facilitate message passing between the two set of nodes, thereby encoding relationship information more effectively. The model is trained using reinforcement learning. Solution are generated in an auto-regressive manner, where each node is sequentially selected based on the current state of the solution.
Diffusion (SL)
We trained the diffusion model to estimate solution matrices using a supervised learning approach, similar to Difusco (Sun and Yang 2023). This method adapts graph-based denoising diffusion models to more naturally formulate combinatorial optimization problems and generate high-quality solutions. By explicitly modeling the node and edge selection process through corresponding random variables, the model effectively captures the problem’s structure. The model is trained through supervised learning.
Diffusion (RL)
Similar to DDPO, as proposed by  (Black et al. 2023), which demonstrates that framing the denoising process as a multi-step decision-making problem enables policy gradient algorithms to directly optimize diffusion models for downstream objectives, we trained our diffusion model within the reinforcement learning framework. We used the REINFORCE algorithm (Williams 1992) to map the denoising process to the MDP framework.
C.3 Training Detail
Genetic algorithm (GA)
Following the implementation of (Kwon et al. 2021), we utilize 25 chromosomes with a mutation rate and crossover ratio both set to 0.3. Among the 25 initial chromosomes, one is initialized with the solution from the SJF heuristic. The best-performing chromosome is retained across all iterations. We run 1000 iterations per instance.
Particle swam optimization (PSO)
Following the implementation of  (Kwon et al. 2021), we utilize 25 particles with an inertial weight of 0.7. Both the cognitive and social constants are set to 1.5. Additionally, one particle is initialized with the solution from the SJF heuristic. We run 1000 iterations per instance.
MatNet
We use the same hyperparameters reported in  (Kwon et al. 2021), with the exception that the number of stages is set to 1 instead of 3.
Diffusion (SL)
Using CP-SAT, we generate 128,000 training samples for supervised learning. Since the solution matrices for PMSP are not square matrices, we use IC/DC’s encoder in combination with Difusco’s decoder, employing 12 layers for both encoder and the decoder. The denoising timestep is set to 1000 during training but reduced to 20 for faster inference. All other hyperparameters, follow those reported in  (Sun and Yang 2023).
Diffusion (RL)
We train our diffusion model with the objective function
$\nabla_{\theta}\mathcal{J}_{\text{DDRL}}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(X_{t-1}|X_{t},c)r(X_{0},c)\right]$ proposed by (Black et al. 2023). We use the same encoder and decoder architecture as IC/DC, consisting of 5 encoder layers and 2 decoder layers. Apart from setting the denoising timestep to 20, all other hyperparameters are consistent with those used in IC/DC (ours).
IC/DC
For PMSP-20 instances, we use 3 layers for both the encoder and the decoder, with the denoising timestep set to 10. For PMSP-50 instances, we utilize 5 encoder layers and 3 decoder layers, with the denoising timestep set to 15. The hyperparameters $lambda_{1}$ and $\lambda_{2}$ are set to 1e-3 and 1e-6, respectively. The IMPROVEMENT step is executed every 30 epochs. Our model is trained using the Adam optimizer, with a batch size of 512 and a learning rate of 4e-4.
Our code is available at:
https://anonymous.4open.science/r/ICDC_opti-8347
Appendix D ATSP definition and baselines
D.1 Tmat class
While random distance matrices can be generated by choosing random integers, such matrices lack meaningful correlations between distances and doesn’t reflect practical scenarios. Instead we are interested in problems which ATSP instances have the triangle inequality so called ”Tmat class” (Kwon et al. 2021; Cirasella et al. 2001). That is, for a distance matrix $D$ with elements $d_{ij}$ representing the distance between city $c_{i}$ and $c_{j}$, if $d(c_{i},c_{j})\geq d(c_{i},c_{k})+d(c_{k},c_{j})$ then we set $d(c_{i},c_{j})=d(c_{i},c_{j})+d(c_{k},c_{j})$, while diagonal elements are maintained as $d(c_{i},c_{i})=0$. We repeat this procedure until no more changes can be made.
D.2 Baseline
Mixed-integer programming
Mixed integer programming (MIP) is an optimization technique used to solve problems where some of the variables are required to be integers while others can be continuous. Solution methods such as branch and bound, branch and cut etc. are used to solve these kind of problems. To we use CPLEX (IBM 2022; Bliek1ú, Bonami, and Lodi 2014), one of the popular commercial optimization software use by the OR community and solve our test instances through benders decomposition (Rahmaniani et al. 2017). The MIP model serves as the mathematical representation of the problem. For ATSP, our MIP model is based on the formulation presented by Miller, Tucker, and Zemlin (1960).
$\text{minimize}\quad\sum^{N}_{i=1}\sum^{N}_{j=1}d(c_{ij}x_{ij})$ (14)
s.t. $\sum^{N}_{i=1}x_{ij}=1\quad\quad j=1,2,3,\cdots,N$ (15)
$\sum^{N}_{j=1}x_{ij}=1\quad\quad i=1,2,3,\cdots,N$ (16)
$u_{i}-u_{j}+(n-1)\cdot x_{ij}\leq n-2\quad i,j=2,\cdots,N$ (17)
$i,j:\text{City index}$
$N:\text{Number of Cities}$
$c_{ij}:\text{Distance from city }i\text{ to city }j$
$x_{ij}:\begin{cases}1&\text{if you move from city $i$ to city $j$}\\ 0&\text{otherwise}\end{cases}.$
$u_{i}:\text{arbitrary numbers representing the order of city $i$in the tour}$
The constraints (15) and  (16) ensure that each city is visited exactly once. Constraint (17) prevents subtours, ensuring that all cities are included in a single tours of length $n$.
Heuristics
As the name suggests, Nearest Neighbor (NN), Nearest Insertion (NI), and Furthest Insertion (FI) are straightforward simple greedy-selection algorithms frequently used as baselines for TSP algorithms. We use the implementations provided by  (Kwon et al. 2021) which are implemented in C++.
LKH3 is a widely recognized state-of-the-art algorithm for addressing constrained TSP and Vehicle Routing Problems (VRP). It employs a local search approach utilizing $k$-opt operations to enhance its solutions. For solving the ATSP instances, we utilize version 3.0.6.
D.3 Training Detail
MatNet
We utilize the checkpoints provided by. (Kwon et al. 2021) and evaluate them on the same problem instances.
Diffusion (SL)
Using LKH-3, we generate 128,000 training samples for supervised learning. We use Difusco’s encoder and decoder, employing 12 layers for both encoder and the decoder. The denoising timestep is set to 1000 during training but reduced to 20 for faster inference. All other hyperparameters, follow those reported in  (Sun and Yang 2023).
Diffusion (RL)
We train our diffusion model with the objective function
$\nabla_{\theta}\mathcal{J}_{\text{DDRL}}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(X_{t-1}|X_{t},c)r(X_{0},c)\right]$ proposed by (Black et al. 2023). We use the same encoder and decoder architecture as IC/DC, consisting of 5 encoder layers and 2 decoder layers. Apart from setting the denoising timestep to 20, all other hyperparameters are consistent with those used in IC/DC.
IC/DC
For ATSP-20 instances, we use 3 layers for both the encoder and the decoder, with the denoising timestep set to 10. For ATSP-50 instances, we utilize 5 encoder layers and 3 decoder layers, with the denoising timestep set to 15. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1e-3 and 1e-6, respectively. The IMPROVEMENT step is executed every 30 epochs. Our model is trained using the Adam optimizer, with a batch size of 256 for 20 node instances and 64 for 50 node instances and a learning rate of 4e-4.
D.4 Complex Feature Data
NP-20
Tour Time $\downarrow$ Gap $\downarrow$ Time $\downarrow$
(Meta-)heuristics Random 3.50 76.43% (0s)
Auto-regressive MatNet (×$\times 256$) 2.60 49.95% (1m)
Ours IC/DC* (×$\times 64$) 1.56 0% (1m)
Table 3: (Results for NP-20). * denotes the baseline for computing the performance gap.
In real-world combinatorial optimization (CO) problems, direct problem data, such as the distance matrix in ATSP or the processing time matrix in PMSP, may not always be explicitly provided. In such cases, solving the problem may require processing a broader range of data. For example, consider a navigation problem (NP) similar to ATSP, where the objective is to minimize the total travel time rather than distance. In this scenario, various types of information that influence travel time must be considered, including the coordinates of each city, time-per-distance data representing the relationships between cities, and traffic information.
We define the actual travel time between $|\mathcal{N}|$ cities as follows:
$[T]^{a}_{i,j}=([R]_{i}-[R]_{j})^{2}\cdot[S]_{i,j}+[F]_{i,j}$
where, for each $i,j\in\mathcal{N}$, $T^{a}\in\mathbb{R}^{|\mathcal{N}|\times|\mathcal{N}|}_{+}$ represents the actual travel time matrix, $R\in\mathbb{R}^{|\mathcal{N}|\times 2}_{+}$ is the coordinate matrix of all cities, $S\in\mathbb{R}^{|\mathcal{N}|\times|\mathcal{N}|}_{+}$ is the reciprocal speed matrix, and $F\in\mathbb{R}^{|\mathcal{N}|\times|\mathcal{N}|}$ is the traffic matrix. The goal is to minimize the total $[T]^{a}_{i,j}$ across the entire route.
In this case, instead of the direct problem data $T^{a}$, we need to consider the complex problem instance $c=(\mathcal{N},R,S,F)$. Typically, methods such as solvers or heuristics require significant expert effort to process such data.
On the other hand, IC/DC is capable of handling these diverse types of data and demonstrates strong generalization performance. By the problem encoder, the information $R$ for each city is processed through an embedding layer with dimension $d$ and then sent to $A$ and $B$. The relational information between cities, $S$ and $F$, is input as $D=S||F\in\mathbb{R}^{|\mathcal{N}|\times|\mathcal{N}|\times 2}$. As shown in Table 3, IC/DC outperforms other baselines and demonstrates strong potential for generalizability. Additionally, an example of the NP is illustrated in Figure 5.
Figure 3: Problem Instance $c$
Figure 4: Solution Visualization
Figure 5: (Left) The instance $c$ is depicted with the coordinate matrix $R$ on the left, the traffic matrix $F$ on the top right, and the reciprocal of speed matrix $S$ on the bottom right. (Right) The figure illustrates the solution for the instance shown on the left. The blue dots represent the coordinates of each city, while the red arrows indicate the travel direction between cities in the solution. The green numbers along the arrows represent the actual travel time between the connected cities.